use laplace transforms to solve ivp

x" + 2x' - 15x = 6delta(t -9), x(0) = -5, x'(0) = 7
Answer #1

Given 11 (D) differential equation, x + 2x- 15o = 66(+9) -5, x(o)=7 Taking Laplace Transform of both sides of the given eq

differential equation Now L ( 675366--] LLos-1) 8 (545) 5+5] (5-3) 5+ 3+ - e B Ij L{F(s)} = f(t) Thun, ({e-ns FCs)} = H (